Celtel is having a competition for it cellphone subscribers.It is giving away $1000 for every 100th caller,a cellphone to every 50th caller and a shopping voucher to every 30th caler.Which caller will receive all three prizes?
Found 2 solutions by stanbon, solver91311:
Answer by stanbon(75887)

Celtel is having a competition for it cellphone subscribers.It is giving away $1000 for every 100th caller,a cellphone to every 50th caller and a shopping voucher to every 30th caler.Which caller will receive all three prizes?
-----
Find the least common multiple of 100, 50, and 30.
-----
100 = 2^2*5*2
50 = 2*5^2
30 = 2*3*5
------
least common multiple = 2^2*5^2*3 = 4*25*3 = 300
-----
The 300th caller will get all three.
Cheers,
Stan H.

Answer by solver91311(24713)



30 is 3 times 10, 50 is 5 times 10, and 1000 is 100 times 10. So 3 times 5 times 100 times 10 = 15,000. Mathematically, the 15000th caller would get all three prizes. Practically speaking, most commercial giveaway programs have fine print rules that say you can only win one prize at a time, hence the real answer is probably nobody would get all three prizes.

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism